# Adaptive Online Bayesian Estimation of Frequency Distributions with Local Differential Privacy

### v1

@@ 4.2.2 Entropy of randomized response
For discrete distributions, entropy measures the uniformity. Hence in the LDP framework, a lower entropy for the randomized response $Y$ implies a more informative $Y$. Based on that observation, a utility function can be defined as the negative entropy of the marginal distribution of $Y$.
@@ 4.2.3 Total variation distance
The TV distance between two discrete probability distributions $\mu,\nu$ on $\mathcal{X}$ is given by
$\textup{TV}(\mu,\nu):=\frac{1}{2}\sum_{x\in\mathcal{X}}|\mu(x)-\nu(x)|.$
@@
relative to its prior $\text{Cat}(\theta)$. The expected amount of change can be formulated as the expectation of the TV distance between the prior and posterior distributions with respect to the marginal distribution of $Y$ given $\theta$. Then a utility function can be defined as
@@ 4.2.4 Expected mean squared error
One can also wish to choose $S$ such that the Bayesian estimator of $X$ given $Y$ has the lowest expected squared error. Specifically, given $k\in\{1,\ldots,K\}$ let $e_{k}$ be a $K\times 1$ vector of 00s except that $e_{k}=1$. A utility function can be defined based on that as
@@
where $\mathbb{E}_{\theta}\left[\|e_{X}-\widehat{e_{X}}(Y)\|^{2}\right]$ is the mean squared error for the estimator $\widehat{e_{X}}$ of $e_{X}$ given $Y$ when $X\sim\text{Cat}(\theta)$ and $Y=\texttt{RRRR}(X|S,\epsilon)$, which is known to be minimized when $\widehat{e_{X}}$ is the Bayesian estimator of $e_{X}$. Proposition 3 provides an explicit formula for this utility function. A proof is given in Appendix A.2.
@@ 4.2.5 Probability of honest response
$U_{6}(\theta,S,\epsilon):=\mathbb{P}(Y=X|S,\theta).$ (10)
@@
$\max_{k=0,\ldots,K-1}U_{6}(\theta,S_{k,\theta},\epsilon)=\max_{S\subset\{1,\ldots,K\}}U_{6}(\theta,S,\epsilon).$
@@ 4.2.6 Semi-adaptive approach
We call this approach “semi-adaptive” because, while it still adapts to $\theta$, it uses the fixed parameter $\alpha$ which is not updated during AdOBEst-LDP. As we will see in Section 7, the best $\alpha$ depends on various parameters such as $\epsilon$, $K$, and the degree of evenness in $\theta$.
@@ 4.3 Computational complexity of utility functions
We now provide the computational complexity analysis of the utility metrics presented in Section 4.2.1-4.2.5, and that of the semi-adaptive approach in Section 4.2.6, as a function of $K$. The first row of Table 1 shows the computational complexities of calculating the utility function for a fixed $S$, and the second row shows the complexities of choosing the best $S$ according to (5). To find (5), the utility function generally needs to be calculated $K$ times, which explains the additional $K$ factor in the computational complexities in the second row.
The least demanding utility function is $U_{6}$, that is based on $\mathbb{P}_{\theta}(Y=X)$, whose complexity is $\mathcal{O}(K)$. Moreover, finding the best $S$ can also be done in $\mathcal{O}(K)$ time because one can compute this utility metric for all $k=0,\ldots,K-1$ by starting with $S=\emptyset$ and populating it incrementally. Also note that the semi-adaptive approach does not use a utility metric and finding $k^{\ast}$ can be done in $\mathcal{O}(K)$ time by summing the components of $\theta$ from largest to smallest until the cumulative sum exceeds the given $\alpha$ parameter. So, its complexity is $\mathcal{O}(K)$.
@@ 5 Posterior sampling
Step 3 of Algorithm 1 requires sampling from the posterior distribution $\Pi(\cdot|Y_{1:n},S_{1:n})$ of $\theta$ given $Y_{1:n}$ and $S_{1:n}$, where $S_{t}$ is the subset selected at time $t$ to generate $Y_{t}$ from $X_{t}$. Let $\pi(\theta|Y_{1:n},S_{1:n})$ denote the pdf of $\Pi(\cdot|Y_{1:n},S_{1:n})$. Given $Y_{1:n}=y_{1:n}$ and $S_{1:n}=s_{1:n}$, the posterior density can be written as
@@
Note that the right-hand side does not include a transition probability for $S_{t}$’s because the sampling procedure of $S_{t}$ given $Y_{1:t-1}$ and $S_{1:t-1}$ does not depend on $\theta$. Furthermore, we assume that the prior distribution $\eta(\theta)$ is a Dirichlet distribution $\theta\sim\textup{Dir}(\rho_{1},\ldots,\rho_{K})$ with prior hyperparameters $\rho_{k}>0$, for $k=1,\ldots,K$.
Unfortunately, exact sampling from the posterior in (11) is difficult. Therefore, we resort to approximate sampling approaches using MCMC. Below we present two MCMC sampling methods. The first method is based on SGLD, which works with random small-sized subsamples and therefore is computationally scalable in $n$. SGLD is an asymptotically exact MCMC sampling approach, targeting the posterior distribution approximately. The second method is a Gibbs sampler, an exact MCMC method unlike the SGLD. However, the computational cost of the Gibbs sampler for $n$ time steps is $\mathcal{O}(n^{2}K)$, making the Gibbs sampler affordable when $n$ is small to moderate, but too expensive for large $n$. In our numerical experiments, we preferred SGLD due to its computational advantage over the Gibbs sampler.
@@ 5.1 Stochastic gradient Langevin dynamics
SGLD is an asymptotically exact MCMC sampling approach that enables the use of subsamples of size $m\ll t$. A direct application of SGLD to generate samples for $\theta$ from the posterior distribution in (11) is difficult. This is because $\theta$ lives in the probability simplex $\Delta$, which makes the task of keeping the iterates for $\theta$ inside $\Delta$ challenging. We overcome this problem by defining the surrogate variables $\phi_{1},\ldots,\phi_{K}$ with
@@
$\phi^{(j)}=\left|\phi^{(j-1)}+\frac{\gamma}{2}\left(\nabla_{\phi}\ln p(\phi^{(j-1)})+\frac{t}{m}\sum_{i=1}^{m}\nabla_{\phi}\ln p(y_{u_{i}}|\phi^{(j-1)})\right)+\gamma W_{t}\right|,\quad W_{t}\sim\mathcal{N}(0,I_{K}).$ (13)
where $u=(u_{1},\ldots,u_{m})$ is a random subsample of $\{1,\ldots,t\}$. In (13), the ‘new’ prior and likelihood functions are
$p(\phi)=\prod_{k=1}^{K}\text{Gamma}(\phi_{k};\alpha_{i},1),\quad p(y|\phi)=h_{S,\epsilon}(y|\theta(\phi)).$ (14)
The reflection in (13) via taking the component-wise absolute value is necessary because each $\phi_{i}^{(j)}$ must be positive. Step 3 of Algorithm 1 can be approximated by running SGLD for some $M>0$ iterations. To exploit the SGLD updates from the previous time, one should start the updates at time $t$ by setting the initial value for $\phi$ to the last SGLD iterate at time $t-1$.
@@ 5.2 Gibbs sampling
An alternative to SGLD is the Gibbs sampler, which operates on the joint posterior distribution of $\theta$ and $X_{1:n}$ given $Y_{1:n}=y_{1:n}$ and $S_{1:t}=s_{1:n}$,
@@
where $p_{s_{t},\epsilon}(x_{t}|y_{t},\theta)$ is defined in (8). Therefore, (15) is a product of $n$ categorical distributions each with support $\{1,\ldots,K\}$. Furthermore, the full conditional distribution of $\theta$ is a Dirichlet distribution due to the conjugacy between the categorical and the Dirichlet distributions. Specifically,
@@ 6 Theoretical analysis
We address two questions concerning AdOBEst-LDP in Algorithm 1 when run with RRRR. (i) Does the targeted posterior distribution based on the observations generated by Algorithm 1 converge to the true value $\theta^{\ast}$? (ii) How frequently does Algorithm 1 with RRRR select the optimum subset $S$ according to the chosen utility function?
@@ 6.1 Convergence of the posterior distribution
  • Let $Q(\cdot|Y_{1:n},S_{1:n},\Theta_{n-1})$ be the probability distribution corresponding to the posterior sampling process for $\Theta_{n}$. Note that, if perfect posterior sampling were used, we would have $Q(A|Y_{1:n},S_{1:n},\Theta_{n-1})=\Pi(A|Y_{1:n},S_{1:n})$; however, when approximate sampling techniques are used to target $\Pi$, such as SGLD or Gibbs sampling, the equality does not hold in general.
@@
    $S^{\ast}_{\theta}:=\{\sigma_{\theta}(1),\ldots,\sigma_{\theta}(k_{\theta}^{\ast})\},\quad k^{\ast}_{\theta}:=\arg\max_{k\in\{0,\ldots,K-1\}}U(\theta,S_{k,\theta},\epsilon)$
@@
where we use the convention that $Q(\mathrm{d}\theta_{0}|Y_{1:0},S_{1:0},\theta_{-1})=\delta_{\theta_{\text{init}}}(\mathrm{d}\theta_{0})$ for an initial value $\theta_{\text{init}}$.
@@ 6.2 Selecting the best subset
Let $S^{\ast}:=S^{\ast}_{\theta^{\ast}}$ be the best subset at $\theta^{\ast}$. In this part, we prove that, if posterior sampling is performed exactly, the best subset is chosen with an expected long-run frequency of $1$. Our result relies on some mild assumptions.
@@
Given any $S\subset\{1,\ldots,K\}$ and $\epsilon>0$, $U(\theta,S,\epsilon)$ is a continuous function of $\theta$ with respect to the $L_{2}$-norm.
@@
Suppose Assumptions 2-4 hold and $\Theta_{t}$s are generated by perfect sampling, that is, $Q(A|Y_{1:t},S_{1:t})=\Pi(A|Y_{1:t},S_{1:t})$ for all measurable $A\subseteq\Delta$. Then,
$P_{\theta^{\ast}}(S_{t}=S^{\ast})\rightarrow 1.$ (18)
@@ 7 Numerical results
We tested the performance of AdOBEst-LDP run with RRRR for randomization when the subset $S$ in RRRR is determined according to a utility function in Section 4.2.111The MATLAB code at https://github.com/soneraydin/AdOBEst_LDP can be used to reproduce the results obtained in this paper. We compared AdOBEst-LDP when combined with each of the utility functions defined in Sections 4.2.1-4.2.5 with its non-adaptive counterpart when SRR is used to generate $Y_{t}$ at all steps. We also included the semi-adaptive subset selection method in Section 4.2.6 into the comparison. For the semi-adaptive approach, we obtained results for 5 different values of its $\alpha$ parameter, namely $\alpha\in\{0.2,0.6,0.8,0.9,0.95\}$.
@@
Finally, the comparison among the methods was repeated for all the combinations ($K\times\epsilon\times\epsilon_{1}\times\rho$) of $K\in\{10,20\}$, $\epsilon\in\{0.5,1,5\}$, $\epsilon_{1}\in\{0.8\epsilon,0.9\epsilon\}$, and $\rho\in\{0.01,0.1,1\}$.
The accuracy results for the methods in comparison are summarized in Figures 3 and 4 in terms of the error given in (20). The box plots are centered at the error median and the whiskers stretch from the minimum to the maximum over the 50 MC runs excluding the outliers. When the medians are compared, the fully adaptive algorithms, which use a utility function to select $S_{t}$, yield comparable results to the best semi-adaptive approach in both figures. As one may expect, the non-adaptive approach yielded the worst results in general, especially in the high-privacy regimes (smaller $\epsilon$) and uneven $\theta^{\ast}$ (smaller $\rho$). We also observe that, while most utility metrics are generally robust, the one based on FIM seems sensitive to the choice of $\epsilon_{1}$ parameter. This can be attributed to the fact that the FIM approaches singularity when $\epsilon_{2}$ is too small, which is the case if $\epsilon_{1}$ is chosen too close to $\epsilon$. Supporting this, we see that when $\epsilon_{1}=0.8\epsilon$, the utility metric based on FIM becomes more robust. Another remarkable observation is that the utility function based on the probability of honest response, $U_{6}$, has competitive performance despite being the lightest utility metric in computational complexity. Finally, while the semi-adaptive approach is computationally less demanding than most fully adaptive versions, the results show it can dramatically fail if its $\alpha$ hyperparameter is not tuned properly. In contrast, the fully adaptive approaches adapt well to $\epsilon$ or $\rho$ and do not need additional tuning.
In addition to the error graphs, the heat maps in Figures 5 and 6 show the effect of parameters $\rho$ and $\epsilon$ on the average cardinality of the subsets $S$ chosen by each algorithm (again, averaged over 50 Monte Carlo runs). According to these figures, increasing the value of $\rho$ causes an increase in the cardinalities of subsets chosen by each algorithm (except the nonadaptive one, since it uses all $K$ categories, rather than a smaller subset). This is expected, since higher $\rho$ values cause $\text{Cat}(\theta^{\ast})$ to be closer to the uniform distribution, thus causing $X$ to be more evenly distributed among the categories. Moreover, for small $\rho$, increasing the value of $\epsilon$ causes a decrease in the cardinalities of these subsets, which can be attributed to a higher $\epsilon$ leading to a more accurate estimation. When we compare the utility functions for the adaptive approach among themselves, we observe that for $\epsilon_{1}=0.8\epsilon$, the third utility function (TV1) uses the subsets with the largest cardinality (on average). However, when we increase the $\epsilon_{1}$ value to $\epsilon_{1}=0.9\epsilon$, the second utility function (FIM) uses the subsets with the largest cardinality. This might be due to the sensitivity of the FIM-based utility function to the choice of $\epsilon_{1}$ parameter that we mentioned before, which affects the invertibility of the Fisher information matrix when $\epsilon_{1}$ is too close to $\epsilon$.

### v2

@@ 4.3.2 Entropy of randomized response
For discrete distributions, entropy measures uniformity. Hence, in the LDP framework, a lower entropy for the randomized response $Y$ implies a more informative $Y$. Based on that observation, a utility function can be defined as the negative entropy of the marginal distribution of $Y$,
@@ 4.3.3 Total variation distance
The TV distance between two discrete probability distributions $\mu,\nu$ on $[K]$ is given by
$\textup{TV}(\mu,\nu):=\frac{1}{2}\sum_{k=1}^{K}|\mu(x)-\nu(x)|.$
@@
relative to its prior $\text{Cat}(\theta)$. The expected amount of change can be formulated as the expectation of the TV distance between the prior and posterior distributions with respect to the marginal distribution of $Y$ given $\theta$. Then, a utility function can be defined as
@@ 4.3.4 Expected mean squared error
One can also wish to choose $S$ such that the Bayesian estimator of $X$ given $Y$ has the lowest expected squared error. Specifically, given $k\in[K]$ let $e_{k}$ be a $K\times 1$ vector of 00s except that $e_{k}=1$. A utility function can be defined based on that as
@@
where $\mathbb{E}_{\theta}\left[\|e_{X}-\widehat{e_{X}}(Y)\|^{2}\right]$ is the mean squared error for the estimator $\widehat{e_{X}}$ of $e_{X}$ given $Y$ when $X\sim\text{Cat}(\theta)$ and $Y=\texttt{RRRR}(X;S,\epsilon)$, which is known to be minimized when $\widehat{e_{X}}$ is the Bayesian estimator of $e_{X}$. Proposition 3 provides an explicit formula for this utility function. A proof is given in Appendix A.2.
@@ 4.3.5 Probability of honest response
$U_{6}(\theta,S,\epsilon):=\mathbb{P}_{\theta}(Y=X|S).$ (11)
@@
$\max_{k=0,\ldots,K-1}U_{6}(\theta,S_{k,\theta},\epsilon)=\max_{S\subset[K]}U_{6}(\theta,S,\epsilon).$
@@ 4.3.6 Semi-adaptive approach
We call this approach “semi-adaptive” because, while it still adapts to $\theta$, it uses the fixed parameter $\alpha$. As we will see in Section 7, the best $\alpha$ depends on various parameters such as $\epsilon$, $K$, and the degree of evenness in $\theta$.
@@ 4.4 Computational complexity of utility functions
We now provide the computational complexity analysis of the utility metrics presented in Section 4.3.1-4.3.5, and that of the semi-adaptive approach in Section 4.3.6, as a function of $K$. The first row of Table 1 shows the computational complexities of calculating the utility function for a fixed $S$, and the second row shows the complexities of choosing the best $S$ according to (6). To find (6), the utility function generally needs to be calculated $K$ times, which explains the additional $K$ factor in the computational complexities in the second row.
The least demanding utility function is $U_{6}$, that is based on $\mathbb{P}_{\theta}(Y=X)$, whose complexity is $\mathcal{O}(K)$. Moreover, finding the best $S$ can also be done in $\mathcal{O}(K)$ time because one can compute this utility metric for all $k=0,\ldots,K-1$ by starting with $S=\emptyset$ and expanding it incrementally. Also note that the semi-adaptive approach does not use a utility metric and finding $k^{\ast}$ can be done in $\mathcal{O}(K)$ time by summing the components of $\theta$ from largest to smallest until the cumulative sum exceeds the given $\alpha$ parameter. So, its complexity is $\mathcal{O}(K)$.
@@
In practice, one can choose among these utility functions depending on the nature of the application. When the number of categories $K$ or the arrival rate of sensitive data is large, we suggest using $U_{6}$ or a semi-adaptive approach. When $K$ and the arrival rate of the personal data are both small, the more computationally demanding utility functions can also be used.
@@
Example 1 (Numerical illustration).
We close this section with an example that shows the benefit of RRRR and the role of $S$. We consider $\theta$ values such that $\theta_{i}/\theta_{i+1}$ is constant for $i=1,\ldots,K-1$. The ratio $\theta_{i}/\theta_{i+1}$ controls the degree of ‘evenness’ in $\theta$: The smaller ratio indicates a more evenly distributed $\theta$. Note that $\theta$ is already ordered in this example; hence, we consider using $S=\{1,\ldots,k\}$ which has the $k$ most likely values for $X$ according to $\theta$. Also, for a given $\epsilon$, we fix $\epsilon_{1}=0.9\epsilon$ and set $\epsilon_{2}$ according to (3).
Figure 2 shows, for a fixed $\epsilon$ and $K=20$, and various values of $k$, the probability of the randomized response being equal to the sensitive information, i.e., $\mathbb{P}_{\theta}(Y=X)$ vs $\theta_{i}/\theta_{i+1}$ when $S=\{1,\ldots,k\}$ in RRRR. (Recall that this probability corresponds to $U_{6}(\theta,S,\epsilon)$.) Comparing this probability with $e^{\epsilon}/(e^{\epsilon}+K-1)$, the probability obtained with $Y=\text{SRR}(X;[K],\epsilon)$, it can be observed that RRRR can do significantly better than SRR if $k$ can be chosen suitably. The plots demonstrate that the “suitable” $k$ depends on $\theta$: While the best $k$ tends to be larger for more even $\theta$, small $k$ becomes the better choice for non-even $\theta$ (large $\theta_{i}/\theta_{i+1}$). This is because, when $\theta_{i}/\theta_{i+1}$ is large, the probability is concentrated on just a few components, and $S$ with a small $k$ captures most of the probability. Moreover, the plots for $\epsilon=1$ and $\epsilon=5$ also show the effect of the level of privacy. In more challenging scenarios where $\epsilon$ is smaller, the gain obtained by RRRR compared to SRR is bigger.
Refer to caption Refer to caption
Figure 2: $\mathbb{P}_{\theta}(Y=X)$ vs $\theta_{i}/\theta_{i+1}$ for all $i=1,\ldots,K-1$ with $K=20$. Left: $\epsilon=1$, Right: $\epsilon=5$.
@@ 5 Posterior sampling
Steps 1-2 of AdOBEst-LDP in Algorithm 1 were detailed in the previous section. In this section, we provide the details of Step 3.
Step 3 of AdOBEst-LDP requires sampling from the posterior distribution $\Pi(\cdot|Y_{1:n},S_{1:n})$ of $\theta$ given $Y_{1:n}$ and $S_{1:n}$ for $n\geq 1$, where $S_{t}$ is the subset selected at time $t$ to generate $Y_{t}$ from $X_{t}$. Let $\pi(\theta|Y_{1:n},S_{1:n})$ denote the pdf of $\Pi(\cdot|Y_{1:n},S_{1:n})$. Given $Y_{1:n}=y_{1:n}$ and $S_{1:n}=s_{1:n}$, the posterior density can be written as
@@
Note that the right-hand side does not include a transition probability for $S_{t}$’s because the sampling procedure of $S_{t}$ given $Y_{1:t-1}$ and $S_{1:t-1}$ does not depend on $\theta^{\ast}$. Furthermore, we assume that the prior distribution $\eta(\theta)$ is a Dirichlet distribution $\theta\sim\textup{Dir}(\rho_{1},\ldots,\rho_{K})$ with prior hyperparameters $\rho_{k}>0$, for $k=1,\ldots,K$.
Unfortunately, the posterior distribution in (12) is intractable. Therefore, we resort to approximate sampling approaches using MCMC. Below, we present two MCMC methods, namely SGLD and Gibbs sampling.
@@ 5.1 Stochastic gradient Langevin dynamics
SGLD is an asymptotically exact gradient-based MCMC sampling approach that enables the use of subsamples of size $m\ll t$. A direct application of SGLD to generate samples for $\theta$ from the posterior distribution in (12) is difficult. This is because $\theta$ lives in the probability simplex $\Delta$, which makes the task of keeping the iterates for $\theta$ inside $\Delta$ challenging. We overcome this problem by defining the surrogate variables $\phi_{1},\ldots,\phi_{K}$ with
@@
$\phi^{(j)}=\left|\phi^{(j-1)}+\frac{\gamma_{n}}{2}\left(\nabla_{\phi}\ln p(\phi^{(j-1)})+\frac{n}{m}\sum_{i=1}^{m}\nabla_{\phi}\ln p_{s_{u_{i}},\epsilon}(y_{u_{i}}|\phi^{(j-1)})\right)+\gamma_{n}W_{j}\right|,\quad W_{j}\sim\mathcal{N}(0,I_{K}).$ (14)
where $u=(u_{1},\ldots,u_{m})$ is a random subsample of $\{1,\ldots,n\}$. In (14), the ‘new’ prior and likelihood functions are
$p(\phi):=\prod_{k=1}^{K}\text{Gamma}(\phi_{k};\alpha_{i},1),\quad p_{s,\epsilon}(y|\phi):=h_{s,\epsilon}(y|\theta(\phi)).$ (15)
The reflection in (14) via taking the component-wise absolute value is necessary because each $\phi_{k}^{(j)}$ must be positive. Step 3 of Algorithm 1 can be approximated by running SGLD for some $M>0$ iterations. To exploit the SGLD updates from the previous time, one should start the updates at time $n$ by setting the initial value for $\phi$ to the last SGLD iterate at time $n-1$.
@@ 5.2 Gibbs sampling
An alternative to SGLD is the Gibbs sampler, which operates on the joint posterior distribution of $\theta$ and $X_{1:n}$ given $Y_{1:n}=y_{1:n}$ and $S_{1:n}=s_{1:n}$,
@@
where $p_{s_{t},\epsilon}(x_{t}|y_{t},\theta)$ is defined in (9). Therefore, (16) is a product of $n$ categorical distributions, each with support $[K]$. Furthermore, the full conditional distribution of $\theta$ is a Dirichlet distribution due to the conjugacy between the categorical and the Dirichlet distributions. Specifically,
@@ 6 Theoretical analysis
We address two questions concerning AdOBEst-LDP in Algorithm 1 when it is run with RRRR whose subset is selected as described in Section 4.3. (i) Does the targeted posterior distribution based on the observations generated by Algorithm 1 converge to the true value $\theta^{\ast}$? (ii) How frequently does Algorithm 1 with RRRR select the optimum subset $S$ according to the chosen utility function?
@@ 6.1 Convergence of the posterior distribution
  • Let $Q(\cdot|Y_{1:n},S_{1:n},\Theta_{n-1})$ be the probability distribution corresponding to the posterior sampling process for $\Theta_{n}$. Note that if exact posterior sampling were used, we would have $Q(A|Y_{1:n},S_{1:n},\Theta_{n-1})=\Pi(A|Y_{1:n},S_{1:n})$; however, when approximate sampling techniques are used to target $\Pi$, such as SGLD or Gibbs sampling, the equality does not hold in general.
@@
    $S^{\ast}_{\theta}:=\{\sigma_{\theta}(1),\ldots,\sigma_{\theta}(k_{\theta}^{\ast})\},\quad\text{with}\quad k^{\ast}_{\theta}:=\arg\max_{k\in\{0,\ldots,K-1\}}U(\theta,S_{k,\theta},\epsilon),$
@@
where we use the convention that $Q(\mathrm{d}\theta_{0}|Y_{1:0},S_{1:0},\theta_{-1})=\delta_{\theta_{\text{init}}}(\mathrm{d}\theta_{0})$ for an initial value $\theta_{\text{init}}\in\Delta$.
@@ 6.2 Selecting the best subset
Let $S^{\ast}:=S^{\ast}_{\theta^{\ast}}$ be the best subset at $\theta^{\ast}$. In this part, we prove that if posterior sampling is performed exactly, the best subset is chosen with an expected long-run frequency of $1$. Our result relies on some mild assumptions.
@@
Given any $S\subset[K]$ and $\epsilon>0$, $U(\theta,S,\epsilon)$ is a continuous function of $\theta$ with respect to the $L_{2}$-norm.
@@
Suppose Assumptions 1-4 hold and $\Theta_{t}$s are generated by exact sampling, that is, $Q(A|Y_{1:t},S_{1:t})=\Pi(A|Y_{1:t},S_{1:t})$ for all measurable $A\subseteq\Delta$. Then,
$\lim_{n\rightarrow\infty}P_{\theta^{\ast}}(S_{n}=S^{\ast})\rightarrow 1.$ (19)
@@ 7 Numerical results
We tested [^1] The MATLAB code at https://github.com/soneraydin/AdOBEst_LDP can be used to reproduce the results obtained in this paper. the performance of AdOBEst-LDP when the subset $S$ in RRRR is determined according to a utility function in Section 4.3. We compared AdOBEst-LDP when combined with each of the utility functions defined in Sections 4.3.1-4.3.5 with its non-adaptive counterpart when SRR is used to generate $Y_{t}$ at all steps. We also included the semi-adaptive subset selection method in Section 4.3.6 into the comparison. For the semi-adaptive approach, we obtained results for five different values of its $\alpha$ parameter, namely $\alpha\in\{0.2,0.6,0.8,0.9,0.95\}$.
@@
Finally, the comparison among the methods was repeated for all the combinations ($K,\epsilon,\kappa,\rho$) of $K\in\{10,20\}$, $\epsilon\in\{0.5,1,5\}$, $\kappa\in\{0.8,0.9\}$, and $\rho\in\{0.01,0.1,1\}$.
@@
The accuracy results for the methods in comparison are summarized in Figures 3 and 4 in terms of the error given in (21). The box plots are centered at the error median, and the whiskers stretch from the minimum to the maximum over the 50 MC runs, excluding the outliers. When the medians are compared, the fully adaptive algorithms, which use a utility function to select $S_{t}$, yield comparable results to the best semi-adaptive approach in both figures. As one may expect, the non-adaptive approach yielded the worst results in general, especially in the high-privacy regimes (smaller $\epsilon$) and uneven $\theta^{\ast}$ (smaller $\rho$). We also observe that, while most utility metrics are generally robust, the one based on FIM seems sensitive to the choice of $\epsilon_{1}$ parameter. This can be attributed to the fact that the FIM approaches singularity when $\epsilon_{2}$ is too small, which is the case if $\epsilon_{1}$ is chosen too close to $\epsilon$. Supporting this, we see that when $\epsilon_{1}=0.8\epsilon$, the utility metric based on FIM becomes more robust. Another remarkable observation is that the utility function based on the probability of honest response, $U_{6}$, has competitive performance despite being the lightest utility metric in computational complexity. Finally, while the semi-adaptive approach is computationally less demanding than most fully adaptive versions, the results show it can dramatically fail if its $\alpha$ hyperparameter is not tuned properly. In contrast, the fully adaptive approaches adapt well to $\epsilon$ or $\rho$ and do not need additional tuning.
@@
In addition to the error graphs, the heat maps in Figures 5 and 6 show the effect of parameters $\rho$ and $\epsilon$ on the average cardinality of the subsets $S$ chosen by each algorithm (again, averaged over 50 Monte Carlo runs). According to these figures, increasing the value of $\rho$ causes an increase in the cardinalities of subsets chosen by each algorithm (except the nonadaptive one since it uses all $K$ categories rather than a smaller subset). This is expected since higher $\rho$ values cause $\text{Cat}(\theta^{\ast})$ to be closer to the uniform distribution, thus causing $X$ to be more evenly distributed among the categories. Moreover, for small $\rho$, increasing the value of $\epsilon$ causes a decrease in the cardinalities of these subsets, which can be attributed to a higher $\epsilon$, leading to a more accurate estimation. When we compare the utility functions for the adaptive approach among themselves, we observe that for $\epsilon_{1}=0.8\epsilon$, the third utility function (TV1) uses the subsets with the largest cardinality (on average). However, when we increase the $\epsilon_{1}$ value to $\epsilon_{1}=0.9\epsilon$, the second utility function (FIM) uses the subsets with the largest cardinality. This might be due to the sensitivity of the FIM-based utility function to the choice of $\epsilon_{1}$ parameter that we mentioned before, which affects the invertibility of the Fisher information matrix when $\epsilon_{1}$ is too close to $\epsilon$.